OpenStudy (anonymous):

The equation of line CD is y = 3x − 3. Write an equation of a line perpendicular to line CD in slope-intercept form that contains point (3, 1). y = 3x + 0 y = −3x − 8 y = negative 1 over 3x + 2 y = − 1 over 3x + 0

OpenStudy (xapproachesinfinity):

The perpendicular should be with slope -1 /3 use that with the point they gave you to find the equation

OpenStudy (xapproachesinfinity):

Use point slope form

OpenStudy (triciaal):

CD y = 3x -3 slope = 3 perpendicular lines the product of the slope = -1 slope of line perpendicular to CD = -1/3 line passes through (3, 1) and (x, y) y - 1 = -1/3(x-3) y = -1/3*x -1/3*(-3) + 1 simplify and match with answer choices

OpenStudy (triciaal):

you should get choice c

OpenStudy (xapproachesinfinity):

I was motivating him to do it by himself so he can do easily with any other assignments